OpenStudy (anonymous):

Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find c if a = 15 and b = 20. A) c = 13 B) c = 25 C) c = 35 D) c = 625

OpenStudy (anonymous):

\[c^2 = a^2 + b^2 \] Pythagoras Theroem

OpenStudy (anonymous):

\[ c=\sqrt{ a^2+b^2}= \sqrt{625}=25 \]